# SPADE: Sparsity-Guided Debugging for Deep Neural Networks

### v1

@@ 4.2 Impact of SPADE on neuron visualization
4.2.1 Resolving multifaceted neurons
Feature visualization is an important tool for examining the working pattern of a neural network. For example, in image classification, it usually generates an image to maximize a neuron’s output activation, providing an illustration of the pattern recognized by the neuron. Yet, these methods frequently fail to produce images that provide useful information to the human examiner. As suggested by Ghiasi et al. (2022); Goh et al. (2021); Nguyen et al. (2016), this issue is in part due to the multifaceted nature of many neurons, i.e., each neuron being associated with several concepts. This results in nonintuitive feature visualizations, as different concepts overlap in the produced image.
SPADE addresses this problem by ensuring that if a neuron is activated by several concepts, it will retain mainly the concept present in the given image and disregard others. Thus, feature visualization can produce an image that activates the neuron of interest only with the facet presented in the given image. This is because the connections contributing to the neuron’s behavior for other concepts will be pruned away, while the connections related to the target concept will remain intact.
@@ 4.2.1 Resolving multifaceted neurons
Figure 3: Two-dimensional example to illustrate the effect of SPADE on feature visualization. The feature visualizations (images generated by Olah et al. (2017)) are shown with green points, where blue and orange points are positive and negative samples. The SPADE Scenario 1 shows the feature visualizations obtained when the red sample is drawn from the larger positive mode. Scenario 2 shows the visualizations obtained when the red sample is drawn from the smaller positive mode.
@@
This property is illustrated for a toy example in Figure 3. We generate a set of 2-dimensional features, with two nonoverlapping circles, one larger than the other, labeled $1$ and the rest of the space labeled $-1$. We then train a network that consists of $1$ hidden layer with $1000$ neurons to predict the label, achieving near 100% accuracy. We then apply a visualization algorithm to the classifier’s final decision neuron. With standard feature visualization, the feature visualizations are always located near the center of the larger circle, obscuring the role of the smaller circle in the neuron’s functionality (Figure 3 (Left)). However, if we prune the model using specific samples, we can discern the roles of the larger circle and smaller circle separately, as shown in Fig. 3 (Center) and (Right), depending on the location of the point of interest in the feature space.
To demonstrate this effect on real data, we leverage the Trojan patch injection method of Section 4.1. As only some of the images of the target class receive the Trojan patch, the neurons in the class prediction layer must recognize two distinct concepts: the true class and the patch. Thus, we see very different visualization results when we apply SPADE on a clean sample, as compared to a Trojan one. We demonstrate this for the Albatross class neuron in Figure 1. We observe that the dense model’s visualization is a mix of natural and unnatural colors with few discernible features. Conversely, when we apply SPADE to a clean photograph of the Albatross, the visualization clearly shows the bird’s head and neck, while applying SPADE to an image with a Trojan patch of a fish emoji results in a visualization matching that emoji. We provide further examples in Appendix G.
@@
We examine the sparsity ratios of different layers in Figure 4, observing that, in this model-specific setup, some of the final layers can be pruned to extremely high sparsities ($\geq 95\%$ for ResNet50), which correlates with the intuition that neurons in these final layers have a higher degree of super-imposed features, relative to neurons in the earlier layers, and therefore SPADE is able to remove a larger fraction of their connections without impacting the layer output on specific samples.
@@ 4.2.2 Human Study
For the evaluation, we randomly sampled 100 misclassified samples. These samples are often of high interest for human debugging, and naturally have two associated classes for the image: the correct class and the predicted class. We used Score-CAM (Wang et al., 2019), a method that has been shown to be class-sensitive, to obtain saliency maps, and corresponding image regions, for each of the two classes. To make this decision more meaningful, we only used samples for which the saliency maps of the two classes have no intersection. To measure relevancy, the image patches were always generated from the dense model. For neuron visualization, we used the method of Olah et al. (2017) implemented in the Lucent/Lucid library. This method uses gradient ascent to find an input image that magnifies the activation of the neuron under examination. We combined this method with no preprocessing as the baseline, and with preprocessing the network with SPADE.
We then showed one of the class feature visualizations, the full image, and the image patches corresponding to the two classes to the evaluators, along with options to either select which of the two regions activates the neuron, or to indicate that the visualization did not enable them to do so. Crucially, we did not disclose the class associated with the neuron. In total, there were were a total of 400 possible human tasks: 100 samples, for which one of two class neurons was interpreted, with the neuron visualization created with or without preprocessing with SPADE. The tasks were chosen randomly from this pool; in total, 24 volunteer evaluators performed 746 rating tasks. We describe the human evaluation and process in more detail and screenshots of sample tasks in Appendix H.
@@
Incorrect attribution $\downarrow$ 20.4% 17.8%
Correct attribution $\uparrow$ 56.7% 69.8%
@@
Figure 4: (Left) Results of human evaluation, measuring the ability of the evaluators to use neuron visualizations to attribute a classification decision to one of two image patches. (Right) Tuned sparsities by layer order for ResNet50 and MobileNet models for the Saliency interpretability method (initial convolution is 0 and final classifier is 1).
@@ Results.
The results of the human evaluation are presented in Figure 4 (left). When the network was preprocessed via SPADE, the users were over 10% more likely to choose to make a decision on which of the patches were responsible for the class prediction (87.4% when SPADE was used, versus 77.1% when it was not). In cases in which the human raters did make a decision, the accuracy was 5.3% higher when SPADE was used (79.9% vs. 73.6%), leading to a major 13.1% increase in net correct attributions. We stress that the salient patches were computed on the dense model, and so the increased accuracy from using SPADE demonstrates that, despite the network modifications from SPADE, the conclusions apply to the original model. Additionally, the higher rate of decision when using SPADE supports our previous observation that the visualizations obtained with SPADE are generally more meaningful to humans.
5 Conclusions and future work
@@
We presented a pruning-inspired method, SPADE, which can be used as a network pre-processing step in a human interpretability pipeline to create interpretability tools are tailored to the input being studied. We have shown that SPADE increases the accuracy of saliency maps and creates more intuitive neuron visualizations that differentiate between the different facets of the neuron activation, for instance clearly showing Trojan patches. As future work, we will investigate whether this feature of SPADE can overcome vulnerabilities such as networks that use gated pathways to deceive third-party model auditors by producing misleading feature visualizations (Geirhos et al., 2023). Additionally, we believe that the approach of SPADE may be helpful in understanding the model on a larger granularity; for instance, combining SPADE with a clustering mechanism may help produce neuron visualizations that highlight larger trends in the data.
Acknowledgments
@@
The authors would like to thank Stephen Casper and Tony Wang for their feedback on this work, and Eldar Kurtic and Elias Frantar for their advice on aspects of the project. This research was supported by the Scientific Service Units (SSU) of IST Austria through resources provided by Scientific Computing (SciComp). EI was supported in part by the FWF DK VGSCO, grant agreement number W1260-N35.
@@ Appendix A Descriptions of saliency methods
In this section, we describe more fully the saliency methods paired with SPADE for the experiments in Section 4.1. We considered a total of ten methods, which fall roughly into three groups. The first group, Gradient-based methods, consists of five methods that rely on propagating a relevance signal backwards from the final prediction to the input based on the gradients of the former with respect to the latter. Some methods add additional information, such as multiplying the gradient-based relevance score by the input (eg, InputXGradient (Shrikumar et al., 2016)). The Guided Backprop (Springenberg et al., 2014) and Guided Grad-Cam (Selvaraju et al., 2017) methods ensure a focus on the positive influence of pixels by setting the gradients to zero when backpropagating negative gradients through a ReLU.
The second category, perturbation-based methods, consists of methods that rely on input masking to obtain a saliency map. Finally, a third category, which we call ’Mixed’, uses a combined approach. Please see Table A.1 for a description of all methods used.
@@
Table A.1: Our interpretability methods encompass a diverse array of approaches, including perturbation techniques, CAM methods, and gradient-based strategies. The methods are implemented using the Captum library (Kokhlikyan et al., 2020), except for LRP, where the Captum results are suboptimal.
Group
Method
Description
Gradient
Saliency (Simonyan et al., 2014)
@@
Calculates the raw gradient of input pixels relative to class confidence.
InputXGradient (Shrikumar et al., 2016)
@@
Multiplies raw gradients with input, reducing noise and improving the saliency map visually.
DeepLift (Shrikumar et al., 2017)
@@
Compares neuron activations with a reference activation calculated using a refrence image to assign neuron’s contributions. Similar saliency map as InputXgradient.
Layer-Wise Relevance Propagation (LRP) (Bach et al., 2015)
@@
Propagates relevance scores from the output to input. Each neuron distribute its relevance to the previous layer’s neurons.
Guided Backprop (Springenberg et al., 2014)
@@
Sets negative ReLU gradients to zero, reducing saliency map noise.
Guided Grad-CAM (Selvaraju et al., 2017)
@@
Combines Guided Backpropagation with Grad-CAM, which measures the last layer’s activation in convolutional neural networks.
Perturbation
Lime (Ribeiro et al., 2016)
@@
Mask some regions of input image and fit a linear model that mimic the original model on the masked images to identify regions’ importance with linear model’s weights.
Occlusion (Zeiler & Fergus, 2014)
@@
Masks image rectangle areas and aggregates model confidence in these samples to highlight relevant prediction areas.
Mixed
IntegratedGradients (Sundararajan et al., 2017)
@@
A smooth variant of InputXgradient, calculates gradients connecting samples to a blank baseline. Then obtain a saliency map using these gradients.
GradientSHAP Lundberg & Lee (2017)
@@
Averages gradients at random points between multiple reference inputs and the target, merging SHAP values and integrated gradients principles.
Appendix B Additional results
B.1 CelebA and Food-101 results on ResNet50
We validate our results on the CelebA and Food-101 datasets (Liu et al., 2015; Bossard et al., 2014). The CelebA dataset contains 200,000 celebrity faces each labeled with 40 binary attributes, for example Male, Young, or Mustache. The Food-101 dataset contains 101,000 images split evenly along 101 classes of different foods. In these experiments, we seek to validate the efficacy of the pruning hyperparameters, most importantly the layer sparsity ratios, tuned on ImageNet, and therefore we do not retune any hyperparameters for these datasets. Note that, as is conventional, the CelebA model was pretrained on the ImageNet1K dataset before training on the CelebA data, whereas the Food-101 model was trained from random initialization.
As in Section 4.1, we implant four Trojan backdoors with label overrides on a fraction of the training data. The backdoors and overrides for CelebA are shown in Table C.6. Hyperparameters of Backdooring process are detailed in Section C. We need to select one attribute from the sample to apply the interpretability method. Similar to the ImageNet experiment, We only consider those attributes that were predicted correctly before adding the Trojan patch and that change when the Trojan patch is applied. We then evaluate the saliency maps for one of these changed attributes.
For Food-101, we follow the ImageNet training recipe detailed in Table C.9. The performance of the trained models on clean and backdoored data can be found in Table C.10. For this dataset we used four emoji as Trojan patches, as shown in Table C.7.
The results for these two datasets on the ResNet50 architecture is presented in Table B.2. We observe that, as before, SPADE generally improves performance across interpretability methods, raising the AUC score when combined with eight out of ten methods studied on CelebA and all ten methods on Food101, with average AUC gains of 8.10% and 11.79%, respectively.
Table B.2: ImageNet, ResNet transferability of sparsity ratio over datasets. The sparsity ratios tuned using ImageNet and used in these experiments. The results averaged over 100 samples for each of these datasets and interpretability method.
@@ B.1 CelebA and Food-101 results on ResNet50
Saliency Method CelebA (ImageNet Pretrained) Food101 (Random Initialization) AUC Pointing Game AUC Pointing Game Dense SPADE $\Delta$ Dense SPADE $\Delta$ Dense SPADE $\Delta$ Dense SPADE $\Delta$ Saliency 73.52 92.81 +19.28 50.67 82.0 +31.33 69.13 94.62 +25.49 33.05 94.92 +61.87 InputXGradient 68.26 92.09 +23.84 32.67 69.33 +36.66 66.09 93.48 +27.39 21.19 90.68 +69.49 DeepLift 87.76 91.21 +3.45 68.0 60.0 -8.0 89.41 95.18 +5.77 72.03 87.29 +15.26 LRP 86.82 96.8 +9.98 34.0 60.0 +26.0 87.26 98.64 +11.38 57.63 88.14 +30.51 GuidedBackprop 97.87 96.63 -1.24 84.67 82.67 -2.0 98.26 98.44 +0.18 93.22 88.14 -5.08 GuidedGradCam 88.89 89.13 +0.24 73.33 71.33 -2.0 97.57 97.61 +0.03 93.22 91.53 -1.69 Lime 75.58 62.42 -13.16 55.33 35.33 -20.0 91.76 93.66 +1.9 53.39 54.24 +0.85 Occlusion 65.12 79.27 +14.15 10.0 64.67 +54.67 75.87 91.45 +15.58 61.02 83.9 +22.88 IntegratedGradients 83.01 93.4 +10.39 64.0 70.0 +6.0 80.02 95.11 +15.1 42.37 89.83 +47.46 GradientShap 80.23 94.25 +14.02 59.33 68.67 +9.34 80.05 95.1 +15.05 43.22 91.53 +48.31 Average 80.71 88.80 +8.10 53.20 66.40 +13.20 83.54 95.33 +11.79 57.03 86.02 +28.99
B.2 MobileNet
In this section, we present the results for the ImageNet and CelebA datasets on the MobileNet-V2 architecture. For MobileNet we exclude depthwise covolutions and only prune pointwise convolutions and linear layers. Further, because the behaviour of LRP is only defined for networks with ReLU activations, we exclude LRP from the analysis. Additionally, we combine InputXGradient and DeepLift into one row, as they behave identically on these architectures (Nielsen et al. (2022), Ancona et al. (2019)).
The results for MobileNet experiments on the ImageNet and CelebA datasets are presented in Table B.3. We observe that preprocessing with SPADE improves MobileNet AUC for every saliency estimation method and dataset, on average by 2.90% for ImageNet and 2.99% for CelebA. Pointing game results are neutral on ImageNet with small changes in average score, but positive on CelebA, with an average improvement of 5.41%.

### v2

@@ 4.2 Impact of SPADE on Neuron Visualization
Figure 3: Two-dimensional example to illustrate the effect of SPADE on feature visualization. The feature visualizations (images generated by (Olah et al., 2017)) are shown with green points, where blue and orange points are positive and negative samples. The SPADE Scenario 1 shows the feature visualizations obtained when the red sample is drawn from the larger positive region. Scenario 2 shows the visualizations obtained when the red sample is drawn from the smaller region.
4.2.1 Visualizing Polysemantic Neurons
Feature visualization is an important tool for examining the working pattern of a neural network. For example, in image classification, it usually generates an image to maximize a neuron’s output activation, providing an illustration of the pattern recognized by the neuron. Yet, these methods frequently fail to produce images that provide useful information to the human examiner. As suggested by (Ghiasi et al., 2022; Goh et al., 2021; Nguyen et al., 2016), this issue is in part due to the polysemantic nature of many neurons, i.e., each neuron being associated with several concepts. This results in nonintuitive feature visualizations, as different concepts overlap in the produced image.
SPADE takes a step towards addressing this problem. We conjecture that, in cases where a neuron may be activated by several concepts (such as images of trees of different species in different seasons and geographies), the connections contributing to the neuron’s affinity for concepts not relevant to the target image will be pruned away, while the connections related to the relevant concept will remain intact. Note, however, that SPADE is not designed to show all possible relevant concepts that activate a neuron, nor elucidate a pathway (or circuit) by which the concept was activated. SPADE does, however, highlight the facet of class neurons that is relevant to the input image, a property that we study quantitatively in 4.2.
@@ 4.2.1 Visualizing Polysemantic Neurons
As additional support for the multifacetism disambiguation conjecture, we conduct a toy experiment. As shown in Figure 3, we generate a set of 2-dimensional features, with two nonoverlapping circles, one larger than the other, labeled $1$ and the rest of the space labeled $-1$. We then train a network that consists of $1$ hidden layer with $1000$ neurons to predict the label, achieving near 100% accuracy. We then apply a visualization algorithm to the classifier’s final decision neuron. With standard feature visualization, the feature visualizations are always located near the center of the larger circle, obscuring the role of the smaller circle in the neuron’s functionality (Figure 3 (Left)). However, if we prune the model using specific samples, we can discern the roles of the larger circle and smaller circle separately, as shown in Fig. 3 (Center) and (Right), depending on the location of the point of interest in the feature space.
To demonstrate this effect on real data, we show two examples of using SPADE to produce image-specific class neuron visualization in Figure 2. Specifically, we examine two images that were incorrectly classified into the “lighter, igniter” class. We observe that the dense model’s visualization does not provide a useful explanation for why these images were misclassified. Conversely, when we apply SPADE, we observe that the class neuron visualisation shows matchsticks in the first case, and flames in the second, providing useful clues as to why the classifier produced incorrect labels. We provide further examples of image-specific class neuron visualizations, where SPADE helps disambiguate between clean and emoji-backdoored images classified into the same class, in Appendix J.
@@
For the neuron visualization setup, some of the final layers can be pruned to extremely high sparsities ($\geq 95\%$ for ResNet50), consistent with the intuition that neurons in these final layers have a higher degree of super-imposed features, relative to neurons in the earlier layers, and therefore SPADE is able to remove a larger fraction of their connections without impacting the layer output on specific samples. We present the sparsities of different layers in Appendix H.
@@ 4.2.2 Human Study
For the evaluation, we randomly sampled 100 misclassified samples. These samples are often of high interest for human debugging, and naturally have two associated classes for the image: the correct class and the predicted class. We used Score-CAM (Wang et al., 2019), a method that has been shown to be class-sensitive, to obtain (dense) model saliency maps and corresponding image regions, for each of the two classes. To prevent ambiguity, we only used samples for which the regions of the two classes have no intersection.
For neuron visualization, we used the method of (Olah et al., 2017) implemented in the Lucent/Lucid library. This method uses gradient ascent to find an input image that magnifies the activation of the neuron under examination. We combined this method with no preprocessing as the baseline, and with SPADE preprocessing. We then randomly selected one of the two relevant classes for an image, and presented its feature visualization, the full image, and the relevance regions for both classes, to the evaluators. We asked them to use the visualization to select which of the two possible relevance regions activates the neuron, or to indicate that they could not do so; crucially, we did not disclose the class associated with the neuron.
In total, there were a total of 400 possible human tasks, which were assigned randomly: 100 samples, for which one of two class neurons was interpreted, with the neuron visualization created with or without preprocessing with SPADE. From these, 24 volunteer evaluators performed 746 rating tasks. More details of the evaluation process are provided in Appendix K.
Table 4: Patch attribution human evaluation results. “Overall success” refers to the ability of the evaluator to identify the same image area as that chosen by Score-CAM.
@@
Agree with Score-CAM $\uparrow$ 56.7% 69.8%
Disagree with Score-CAM $\downarrow$ 20.4% 17.8%
@@
Disagree when not undecided $\downarrow$ 26.4% 20.1%
Overall success $\uparrow$ 56.7% 69.8%
@@ Results.
The results of the human evaluation are presented in Table 4. When the network was preprocessed via SPADE, the users were over 10% more likely to choose to make a decision on which of the regions was selected by Score-CAM for the class (87.4% when SPADE was used, versus 77.1% when it was not). In cases in which the human raters did make a decision, they were more likely to agree with ScoreCAM when SPADE was used (79.9% agreement rate) than when it was not (73.6%). Overall, the evaluators were able to identify the image patch that matched Score-CAM 69.8% of the time when SPADE was used, and 56.7% of the time when it was not. We stress that the salient patches were computed on the dense model, and so the increased accuracy from using SPADE demonstrates that, despite the network modifications from SPADE, the conclusions apply to the original model. Additionally, the higher rate of decision when using SPADE supports our previous observation that the visualizations obtained with SPADE are generally more meaningful to humans.
5 Conclusions, Limitations, and Future Work
@@
We presented a pruning-inspired method, SPADE, which can be used as a network pre-processing step in a human interpretability pipeline to create interpretations that are tailored to the input being studied. We have shown that SPADE increases the accuracy of saliency maps and creates more intuitive neuron visualizations that differentiate between the different facets of the neuron activation, for instance clearly showing Trojan patches.
We have also demonstrated that SPADE enables the application of global interpretability methods, such as feature visualization, in a local context. Global interpretability methods provide an overall view of the model’s decision-making process, while local interpretability methods focus on explaining model behaviour on a single data point. By bridging the gap between global and local interpretability methods, SPADE, enriches the interpretability toolkit.
Limitations and future work.
Although, for all methods, SPADE improves interpretations on average, it is possible that SPADE favors some categories of specific examples over others, in other words, there may be some systemic bias. This is also true for all interpretation methods, and we hope that more work will be done in the future to measure this effect. Further, additional evidence is needed toward our conjecture that the effectiveness of SPADE is due to resolving neuron polysemanticity, especially as investigating this phenomenon may be fruitful in gaining a better mechanistic understanding of the neural network by examining the masks produced by SPADE. Thus, we leave it as future work to explicitly incorporate SPADE into model-wide debugging efforts such as systematic searches for spurious correlations, or circuit identification in networks. Finally, the tuning of SPADE can be costly. We propose some mitigations for this; however, we acknowledge that it may impede practical adoption of SPADE in some cases. Additionally, the computational overhead of SPADE, may require more careful example selection.
As additional future work, we will investigate whether SPADE can overcome additional known vulnerabilities of interpretability methods, such as networks that use gated pathways to produce misleading feature visualizations (Geirhos et al., 2023). We also note that SPADE opens a promising direction for using data to interpret models on a larger granularity; for instance, combining SPADE with a clustering mechanism may help produce neuron visualizations that highlight larger trends in the data, bringing this line of work closer to mechnanistic interpretability literature. We hope that these directions inspire more data-driven interpretability research in this area.
Impact Statement
The goal of this paper is to advance the field of interpretable Machine Learning by demonstrating the positive effect of model pruning on neural network interpretability. The specific social consequences of our work are tied to the use of ML in general; however, we believe that improving the ability to understand complex models is of positive social value.
Acknowledgements
@@
The authors would like to thank Stephen Casper and Tony Wang for their feedback on this work, and Eldar Kurtic for his advice on aspects of the project. This research was supported by the Scientific Service Units (SSU) of IST Austria through resources provided by Scientific Computing (SciComp). EI was supported in part by the FWF DK VGSCO, grant agreement number W1260-N35.
@@ Appendix A Descriptions of Saliency Methods
Table 1: Our interpretability methods encompass a diverse array of approaches, including perturbation techniques, CAM methods, and gradient-based strategies. The methods are implemented using the Captum library (Kokhlikyan et al., 2020), except for LRP, for which we use (Nam et al., 2019).
Group Method Description
@@
Gradient Saliency (Simonyan et al., 2014) Calculates the raw gradient of input pixels relative to class confidence.
@@
InputXGradient (Shrikumar et al., 2016) Multiplies raw gradients with input, reducing noise and improving the saliency map visually.
@@
DeepLift (Shrikumar et al., 2017) Compares neuron activations with a reference activation calculated using a reference image to assign neuron’s contributions. Similar saliency map as InputXgradient.
@@
Layer-Wise Relevance Propagation (LRP) (Bach et al., 2015) Propagates relevance scores from the output to the input. Each neuron distributes its relevance to the previous layer’s neurons.
@@
Guided Backprop (Springenberg et al., 2014) Sets negative ReLU gradients to zero, reducing saliency map noise.
@@
Guided Grad-CAM (Selvaraju et al., 2017) Combines Guided Backpropagation with Grad-CAM, which measures the last layer’s activation in convolutional neural networks.
@@
Perturbation Lime (Ribeiro et al., 2016) Mask some regions of the input image and fit a linear model that mimics the original model on the masked images to identify regions’ importance with the linear model’s weights.
@@
Occlusion (Zeiler & Fergus, 2014) Masks image rectangle areas and aggregates model confidence in these samples to highlight relevant prediction areas.
@@
Mixed IntegratedGradients (Sundararajan et al., 2017a) A smooth variant of InputXgradient, calculates gradients connecting samples to a blank baseline. Then obtain a saliency map using these gradients.
@@
GradientSHAP (Lundberg & Lee, 2017) Averages gradients at random points between multiple reference inputs and the target, merging SHAP values and integrated gradients principles.
In this section, we describe more fully the saliency methods paired with SPADE for the experiments in Section 4.1. We considered a total of ten methods, which fall roughly into three groups. The first group, Gradient-based methods, consists of five methods that rely on propagating a relevance signal backward from the final prediction to the input based on the gradients of the former with respect to the latter. Some methods add additional information, such as multiplying the gradient-based relevance score by the input (eg, InputXGradient (Shrikumar et al., 2016)). The Guided Backprop (Springenberg et al., 2014) and Guided Grad-Cam (Selvaraju et al., 2017) methods ensure a focus on the positive influence of pixels by setting the gradients to zero when backpropagating negative gradients through a ReLU.
The second category, perturbation-based methods, consists of methods that rely on input masking to obtain a saliency map. Finally, a third category, which we call ’Mixed’, uses a combined approach. Please see Table 1 for a description of all methods used.
Appendix B Layer Sparsity Tuning
In this section, we discuss alternative approaches to tuning layer sparsities. As discussed in Section 4.1, we obtain our best results by inserting Trojan patches into the model, which are then used to tune sparsity ratios. We demonstrated that this performs well on the Insertion/Deletion metrics, even when the samples passed through the model are clean. Here, we additionally explore a rule-of-thumb pattern, where target sparsity ratios are chosen to increase linearly from 0 sparsity in the initial convolution to 99% sparsity in the final FC layer. For convenience, rather than using these exact sparsity ratios, we choose the closest sparsity ratio from the ones used in other experiments (0, 20%, 40%, 60%, 80%, 90%, 95%, 99%).
We observe that while using tuned sparsities is more effective than the linear schedule described above, even this simplified version substantially improves over the baseline version, adding an average of 3.42% AUC on the Trojan patch discovery task.
We note that while this simple rule works well with the SPADE/OBC method, we have not found it to work well with the FastSPADE method, likely because a less accurate pruner requires more precise target setting. We also found that target sparsities obtained with the SparseGPT pruner transfer well to OBC, though the converse is not true.
Table 2: ResNet50 results on the ImageNet dataset, averaged over 140 samples. ”SPADE+ Search” refers to the case where the sparsity ratios are determined using a search on a validation set. ”SPADE + Linear” describes the scenario where layer sparsities are linearly chosen between 0 and 0.99, with the input layer assigned a 0 sparsity ratio.
@@ Appendix B Layer Sparsity Tuning
Saliency Method Dense SPADE+Search SPADE+Linear Saliency 86.92 95.32 91.58 InputXGradient 83.77 93.73 88.77 DeepLift 93.47 95.85 94.99 LRP 90.05 99.11 98.15 GuidedBackprop 95.22 96.45 95.59 GuidedGradCam 97.82 98.12 97.87 Lime 91.93 95.84 94.34 Occlusion 86.09 93.73 89.27 Integrated Gradients 87.86 94.77 92.34 GradientSHAP 87.74 94.85 92.15 Average 90.09 95.78 93.51
B.1 Layer Sparsity Search with Smaller Number of Examples
We additionally experiment with using a smaller number of examples to tune the sparsity ratios; the number of examples used has a linear effect on the time to tune. Therefore, we tuned the FastSPADE method similarly to our Trojan identification experiments in Section 4.1, but only using 30 examples per method. The results are shown in Table 3. We observe that, while the average accuracy drops slightly as compared to using more examples for tuning, the quick-tuned method still outperforms the baselines of SparseFC method of Wong et al. (2021) as well as using the dense model without any preprocessing.
Table 3: Trojan patch AUC for FastSPADE calibrated on 30 samples per method, compared to regular FastSPADE, the dense model, and to the Sparse FC method of Wong et al. (2021).
Saliency Method FastSPADE (30 ex.) FastSPADE (100 ex.) SparseFC Dense
Saliency 92.53 93.91 88.05 87.87
InputXGradient 90.99 90.61 85.59 85.44
LRP 97.81 98.03 93.99 90.81
GuidedGradCam 97.4 97.75 98 98.03
DeepLift 95.33 95.07 94.21 94.1
Gradient SHAP 93.23 93.8 89.82 89.51
Occlusion 91.18 90.9 87.84 88.29
Lime 92.44 93.94 91.83 90.69
GuidedBackprop 95.74 95.81 95.82 95.73
IntegratedGradients 93.08 93.55 89.89 89.61
Average 93.97 94.34 91.5 91.01
B.2 Sparsity Ratio Search with FastSPADE
We additionally experimented with using FastSPADE to tune the layer ratios for use with SPADE. This can be advantageous, as FastSPADE is faster to execute. We show the results in Table 4. We observe that, while SPADE tuning slightly outperforms FastSPADE, both show substantial improvement over omitting SPADE, as well as over the SparseFC method of (Wong et al., 2021).
Table 4: Trojan patch AUC for SPADE calibrated using FastSPADE, compared to regular SPADE, the dense model, and to the Sparse FC method of Wong et al. (2021).
Saliency Method SPADE (FastSPADE tuning) SPADE SparseFC Dense
Saliency 93.81 96.21 88.05 87.87
InputXGradient 92.3 95.1 85.59 85.44
LRP 98.14 99.21 93.99 90.81
GuidedGradCam 97.66 98.37 98 98.03
DeepLift 95.25 96.55 94.21 94.1
Gradient SHAP 93.83 96.03 89.82 89.51
Occlusion 90.84 95.4 87.84 88.29
Lime 93.67 95.47 91.83 90.69
GuidedBackprop 95.79 97.08 95.82 95.73
IntegratedGradients 94.54 96.1 89.89 89.61
Average 94.58 96.55 91.5 91.01
B.3 Transferability of Layer Sparsity Targets across Datasets
We validate the transferability of layer sparsity tunings obtained on ImageNet on the CelebA and Food-101 datasets (Liu et al., 2015; Bossard et al., 2014). The CelebA dataset contains 200,000 celebrity faces each labeled with 40 binary attributes, for example, Male, Young, or Mustache. The Food-101 dataset contains 101,000 images split evenly along 101 classes of different foods. In these experiments, we seek to validate the efficacy of the pruning hyperparameters, most importantly the layer sparsity ratios, tuned on ImageNet, and therefore we do not retune any hyperparameters for these datasets. Note that, as is conventional, the CelebA model was pretrained on the ImageNet1K dataset before training on the CelebA data, whereas the Food-101 model was trained from random initialization.
As in Section 4.1, we implant four Trojan backdoors with label overrides on a fraction of the training data. The backdoors and overrides for CelebA are shown in Table 12. Hyperparameters of the Backdooring process are detailed in Appendix F. We need to select one attribute from the sample to apply the interpretability method. Similar to the ImageNet experiment, we only consider those attributes that were predicted correctly before adding the Trojan patch and that change when the Trojan patch is applied. We then evaluate the saliency maps for one of these changed attributes.
For Food-101, we follow the ImageNet training recipe detailed in Table 15. The performance of the trained models on clean and backdoored data can be found in Table 16. For this dataset, we used four emoji as Trojan patches, as shown in Table 13.
The results for these two datasets on the ResNet50 architecture are presented in Table 5. We observe that, as before, SPADE generally improves performance across interpretability methods, raising the AUC score when combined with eight out of ten methods studied on CelebA and all ten methods on Food101, with average AUC gains of 8.10% and 11.79%, respectively.